### v1

@@ 5 The Charging Scheme
We proceed similarly. We combine the demand states when we go up the tree $T$, by constructing a series of flow problems (i.e., demand matrices) to update one demand state to the next (recall Definition 3.18). Instead of showing that these matrices can be routed with low congestion, we charge the edges of the cut $(B^{\prime},W^{\prime})$ to cover the demands that these demand matrices induce across the cut (so that the charge incurred at each step is at least the demand of the demand matrix across the cut). This is made explicit below.
@@
Initially, each edge of $(B^{\prime},W^{\prime})$ is charged 0. We convert the given valid demand state $P$ to multiple demand states $P_{\{v\}^{\prime}}$, one for each vertex $v\in V$. For each $v\in V$, its demand vector $P(v)$ is evenly distributed among the split nodes $x_{e}$, incident to $v$. Formally, each split node $x_{e}$ receives the demand vector $P_{\{v\}^{\prime}}(x_{e})=\frac{1}{\deg(v)}P(v)$. Observe that $\left\|P(v)\right\|_{1}=\operatorname{dem}_{P}{(\{v\},V\setminus\{v\})}\leq\operatorname{cap}_{G}(\{v\},V\setminus\{v\})=\deg_{G}(v)$, where the inequality is due to the fact that $\{v\}$ is a cluster in $T$ and that $P$ is $1$-respected by $T$. In particular, each split node $x_{e}$ has a total load of at most $1$: $\left\|P_{\{v\}^{\prime}}(x_{e})\right\|_{1}=\frac{1}{\deg(v)}\left\|P_{\{v\}^{\prime}}\right\|_{1}\leq 1$. Each non-split vertex of $V$ a load of 0.
@@
We combine demand states as we go up the decomposition tree $T$. We consider each cluster in $T$, and replace the demand states $P_{S_{1}^{\prime}},\ldots,P_{S_{r}^{\prime}}$ that are supported on its sub-clusters $S_{1}^{\prime},\ldots,S_{r}^{\prime}$ with a new demand state that is supported on $S^{\prime}$. This is done in Theorem 5.5 below. Note that our demand states are not valid demand states. They will, however, satisfy the following invariant:
@@
Note that $P_{S^{\prime}}$ satisfying the invariant does not necessarily imply that $P_{S^{\prime}}$ is a valid demand state. It can be seen that $P_{\{v\}^{\prime}}$ satisfies the invariant with $\alpha=1$ for the cluster $\{v\}$.
@@
Let $S$ be a cluster, and let $S_{1},\ldots,S_{t}$ be its sub-clusters (i.e., $\{S_{1},\ldots S_{t}\}=\{L_{1},\ldots,L_{\ell}\}\cup\{R_{1},\ldots,R_{r}\}$), computed by Theorems 4.1 and 4.10 with parameter $0<\uptau\leq 1$. For each $1\leq i\leq t$, assume that $P_{S^{\prime}_{i}}$ is a demand state satisfying Invariant 5.4 for $S_{i}$ with load $\alpha\geq 1$. Let $P^{\mathrm{before}}_{S^{\prime}}=\sum_{i=1}^{t}P_{S^{\prime}_{i}}$ be their sum. Then, we can replace $P^{\mathrm{before}}_{S^{\prime}}$ with a demand state $P^{\mathrm{after}}_{S^{\prime}}$ which satisfies the following:
@@
    $G^{\prime}[S^{\prime}]$ $\Omega\!\left(\frac{\uptau}{\alpha\log n}\right)$-respects $P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}$.
@@
Result (2) of Theorem 5.5 means that if we apply a charge on the edges $e\in E(B^{\prime},W^{\prime})\cap G^{\prime}[S^{\prime}]$ (to cover the demand of $P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}$ across the cut $(B^{\prime},W^{\prime})$) as follows:
$\mathrm{charge}_{S}\mathrel{\mathop{:}}=\frac{\operatorname{dem}_{\left(P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}\right)}(B^{\prime},W^{\prime})}{\operatorname{cap}\!\left(E(B^{\prime},W^{\prime})\cap G^{\prime}[S^{\prime}]\right)},$
then we have that $\mathrm{charge}_{S}=O\!\left(\frac{\alpha\log n}{\uptau}\right)$. Note that the charged edges are in the subdivision graph.
In other words, Theorem 5.5 shows that the demand states supported on the sub-clusters $S_{1},\ldots,S_{t}$ can be merged into a single demand state on $S$, while only slightly increasing the load (by a factor of $1+O(\uptau)$) on the boundary edges of $S$. The cost of this operation is that we add $O\left(\frac{\alpha\log n}{\uptau}\right)$ to the charge of each edge $e\in E(B^{\prime},W^{\prime})\cap G^{\prime}[S^{\prime}]$. The proof of this theorem is given in Section 5.1.
@@
We prove that the conditions of Corollary 5.3 are satisfied. Let $P$, $(B,W)$, and $(B^{\prime},W^{\prime})$ be as in the discussion following Corollary 5.3. Set $\uptau=\frac{1}{\log n}$. We process $P$ in a bottom-up manner: first we convert it to a set of demand states $P_{\{v\}^{\prime}}$ as shown above. Then, for each cluster in $T$, we apply Theorem 5.5 and Remark 5.6, accumulating the charges incurred at each application of the theorem. Let $\mathrm{charge}_{S}(e)\mathrel{\mathop{:}}=\mathrm{charge}_{S}$ denote the charge assigned to an edge $e\in E(B^{\prime},W^{\prime})\cap G^{\prime}[S^{\prime}]$ when Theorem 5.5 is applied to cluster $S$.
After applying the theorem for the final cluster $V$, the resulting demand state $P_{V^{\prime}}$ is actually the zero demand state on $V^{\prime}$. This follows from Invariant 5.4(1), since the cluster $V$ has no boundary edges. We now explain why the accumulated charges are enough to cover $\operatorname{dem}_{P}(B,W)$.
Denote by $\mathcal{Q}$ the set of current “active” clusters, such that initially $\mathcal{Q}_{0}=\{\{v\}\mid v\in V\}$. Then in each iteration we remove $S_{1},\ldots,S_{t}$ from $\mathcal{Q}$ and add $S$. Let $P_{\mathcal{Q}}$ be the sum of all demand states $P_{S^{\prime}}$, for $S\in\mathcal{Q}$.
@@
By induction, for each $S\in\mathcal{Q}$, $P_{S^{\prime}}$ satisfies Invariant 5.4. Therefore, for each $k\in\mathcal{K}$,
@@
The following claim, whose proof is deferred to Appendix A, shows that $\operatorname{dem}_{P}(B,W)$ and $\operatorname{dem}_{P_{\mathcal{Q}_{0}}}(B^{\prime},W^{\prime})$ are the same up to some minor error that arises due to split nodes on the cut $(B^{\prime},W^{\prime})$ of $G^{\prime}$.
@@
$\operatorname{dem}_{P}(B,W)\leq\operatorname{dem}_{P_{\mathcal{Q}_{0}}}(B^{\prime},W^{\prime})+\operatorname{cap}_{G}(B,W)$.
Consider a replacement of $\mathcal{Q}$ by $\mathcal{Q}^{\prime}=\mathcal{Q}\cup\{S\}\setminus\{S_{1},S_{2},\ldots,S_{t}\}$ when we apply Theorem 5.5. Recall that $P^{\mathrm{before}}_{S^{\prime}}=\sum_{i=1}^{t}P_{S^{\prime}_{i}}$ is replaced by $P^{\mathrm{after}}_{S^{\prime}}$. Let $P^{\mathrm{before}}_{\mathcal{Q}^{\prime}}\mathrel{\mathop{:}}=P_{\mathcal{Q}}$ be the sum of active demand states before the change and let $P^{\mathrm{after}}_{\mathcal{Q}^{\prime}}\mathrel{\mathop{:}}=P_{\mathcal{Q}}-P^{\mathrm{before}}_{S^{\prime}}+P^{\mathrm{after}}_{S^{\prime}}$ be the sum after the change. We have $\operatorname{dem}_{\left(P^{\mathrm{before}}_{\mathcal{Q}^{\prime}}-P^{\mathrm{after}}_{\mathcal{Q}^{\prime}}\right)}(B^{\prime},W^{\prime})=\operatorname{dem}_{\left(P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}\right)}(B^{\prime},W^{\prime})$. This means that after accumulating all the charges in the applications of Theorem 5.5 on all clusters $S\in T$, the total charge satisfies
$\sum_{e\in E(B^{\prime},W^{\prime})}{\mathrm{charge}(e)}$ $=\sum_{e\in E(B^{\prime},W^{\prime})}\sum_{S\in T}{\mathrm{charge}_{S}(e)}=\sum_{S\in T}\sum_{e\in E(B^{\prime},W^{\prime})\cap G^{\prime}[S^{\prime}]}{\mathrm{charge}_{S}(e)}$
$\geq\sum_{S\in T}{\operatorname{dem}_{\left(P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}\right)}(B^{\prime},W^{\prime})}\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu2\mkern-1.5mu)}}{{\geq}}\operatorname{dem}_{\left(P_{\mathcal{Q}_{0}}-P_{V^{\prime}}\right)}(B^{\prime},W^{\prime})=\operatorname{dem}_{P_{\mathcal{Q}_{0}}}(B^{\prime},W^{\prime})$
$\geq\operatorname{dem}_{P}(B,W)-\operatorname{cap}_{G}(B,W)~,$
where Inequality (2) uses Fact 3.17. By charging an additional unit to each cut edge, we get $\operatorname{dem}_{P}(B,W)\leq\sum_{e\in E(B^{\prime},W^{\prime})}{\mathrm{charge}(e)}$.
To conclude, we bound the total charge on each edge as follows. The height of the decomposition tree is $O(\log n)$ and $\uptau=\frac{1}{\log n}$, which means that $\alpha$ is a constant for each cluster. Therefore, by Theorem 5.5 and Remark 5.6, each edge is charged $O(\log^{2}n)$ for each cluster it is contained in, for a total charge of $O(\log^{3}n)$ per edge, which gives $\operatorname{dem}_{P}(B,W)=O(\log^{3}n)\cdot\operatorname{cap}_{G^{\prime}}(B^{\prime},W^{\prime})=O(\log^{3}n)\cdot\operatorname{cap}_{G}(B,W)$, as we wanted. Thus, the conditions of Corollary 5.3 are satisfied, and Theorem 2.9 follows. ∎
@@
The goal of this section is to prove Theorem 5.5. Recall the notations of this theorem: $S\in T$ is a cluster, and let $S_{1},\ldots,S_{t}$ be its sub-clusters (i.e., $\{S_{1},\ldots S_{t}\}=\{L_{1},\ldots,L_{\ell}\}\cup\{R_{1},\ldots,R_{r}\}$), computed by Theorems 4.1 and 4.10 with parameter $0<\uptau\leq 1$. For $i=1\ldots t$, $P_{S^{\prime}_{i}}$ is a demand state satisfying Invariant 5.4 for $S_{i}$ with a given load $\alpha\geq 1$.
Let $P^{\mathrm{before}}_{S^{\prime}}\mathrel{\mathop{:}}=\sum_{i=1}^{t}P_{S^{\prime}_{i}}$ be the sum of the demand states on the sub-clusters. By Invariant 5.4(1), all the non-zero demands in $P^{\mathrm{before}}_{S^{\prime}}$ are residing at nodes $X_{F}\cup X_{Y}\cup X_{B}$ of $S^{\prime}$ (these are the boundary edges of $\{S_{i}\}_{i=1}^{t}$). Given these demand vectors, we wish to “solve” some of these demands by updating the demand states such that a positive demand and a negative demand of the same commodity are moved to the same vertex, thus canceling out, and then “send” the remaining demands to $X_{B}$ (the boundary of $S$). In other words, we replace $P^{\mathrm{before}}_{S^{\prime}}$ by a demand state $P^{\mathrm{after}}_{S^{\prime}}$, where the demands of $P^{\mathrm{after}}_{S^{\prime}}$ lie on $X_{B}$ (i.e., satisfying Invariant 5.4 for $S$), and we show that $G^{\prime}[S^{\prime}]$ $\Omega(\frac{\uptau}{\alpha\log n})$-respects $P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}$.
This uses a sequence of demand state updates:
@@
    Denote $P^{(1)}_{S^{\prime}}\mathrel{\mathop{:}}=P^{\mathrm{before}}_{S^{\prime}}$. First, we deal with demands that originated inside $L$, i.e., with the demand states $P^{(1)}_{L^{\prime}}=\sum_{i}{P_{L^{\prime}_{i}}}$, where $L=\bigcup_{i}{L_{i}}$. We move the demands of $P^{(1)}_{L^{\prime}}$ that reside on $X_{Y}\cap L^{\prime}$ to $X_{F}$, using the flow guaranteed by Property 2 of Theorem 4.10. This flow is routable in the graph and in particular, by Fact 2.7, respected by the graph.
@@
    Then, we use the all-to-all multi-commodity flow problem on $X_{F}$ to update the demand state and spread the demands uniformly on $X_{F}$. By Theorem 4.1 we know that this problem is respected by the graph. This means that positive and negative demands of the same commodity cancel out, and we are only left with demand that has to enter or leave the cluster $L$. Because $L$ is a node in the hierarchical decomposition tree $T$ and $P$ is respected by $T$, the total amount of leftover demand is at most $\operatorname{cap}_{G}(L,V\setminus L)$.
@@
    Then, we use the reverse flow from step (1) to route the demands back to $X_{Y}$. [^11] In fact, we route to only some edges of $Y$ (see the proof for details).
@@
    Finally, we take all the demands that are currently on $X_{Y}$ (either from $P_{R^{\prime}_{i}}\mathrel{\mathop{:}}=P^{(1)}_{S^{\prime}}-P^{(1)}_{L^{\prime}}$ or the resulting updated demands), and use the flow ensured by Property 1 of Theorem 4.10 to route them to $X_{B}$. This flow is routable in the graph and, by Fact 2.7, respected by the graph.
In all of these steps, we show that $G^{\prime}[S^{\prime}]$ respects the updates, not necessarily bounding the actual congestion that would be incurred when routing these flows. We split the above process into lemmas.
Let $P^{(1)}_{L^{\prime}}\mathrel{\mathop{:}}=\sum_{i=1}^{\ell}P_{L^{\prime}_{i}}$ be the sum of the demand states on the sub-clusters within the core $L$. The following lemma captures steps $1$ and $2$ in the above process.
@@
    $G^{\prime}[S^{\prime}]$ $\Omega\!\left(\frac{\uptau}{\alpha\log n}\right)$-respects $P^{(1)}_{L^{\prime}}-P^{(3)}_{L^{\prime}}$.
@@
Initially, Invariant 5.4(3) gives that every boundary split-node $x_{b}\in X_{B}$ in the graph $G^{\prime}[S^{\prime}]$ has load at most $\alpha$ in $P^{(1)}_{S^{\prime}}$, while vertices of $(X_{F}\cup X_{Y})\setminus X_{B}$ may have load at most $2\alpha$ (each of these vertices is a boundary node of two of the sub-clusters $\{S_{i}\}_{i=1}^{t}$, so two sub-clusters may have placed a load of $\alpha$ on it). That is, $\sum_{k\in\mathcal{K}}{|P^{(1)}_{S^{\prime}}(x_{b},k)|}\leq\alpha$ for boundary nodes $x_{b}\in X_{B}$, and $2\alpha$ for nodes in $(X_{F}\cup X_{Y})\setminus X_{B}$. Due to the invariant, $P^{(1)}_{L^{\prime}}$ is supported only on $L^{\prime}$.

### v2

@@ 5 The Charging Scheme
We proceed similarly. We combine the demand states when we go up the tree $T$, by constructing a series of flow problems (i.e., demand matrices) to update one demand state to the next (recall Definition 3.17). Instead of showing that these matrices can be routed with low congestion, we charge the edges of the cut $(U^{\prime},W^{\prime})$ to cover the demands that these demand matrices induce across the cut. That is, in each such update step, we assign to every cut edge an amount of charge, so that the total assigned charge is at least the demand that the corresponding demand matrix induces across the cut. This is made explicit below.
@@
Initially, each edge of $(U^{\prime},W^{\prime})$ is charged 0. We convert the given valid demand state $P$ to multiple demand states $P_{\{v\}^{\prime}}$, one for each vertex $v\in V$. For each $v\in V$, its demand vector $P(v)$ is evenly distributed among the split nodes $x_{e}$, incident to $v$. Formally, each split node $x_{e}$ receives the demand vector $P_{\{v\}^{\prime}}(x_{e})=\frac{1}{\deg(v)}P(v)$. Observe that $\left\|P(v)\right\|_{1}=\operatorname{dem}_{P}{(\{v\},V\setminus\{v\})}\leq\operatorname{cap}_{G}(\{v\},V\setminus\{v\})=\deg_{G}(v)$, where the inequality is due to the fact that $\{v\}$ is a cluster in $T$ and that $P$ is $1$-respected by $T$. In particular, each split node $x_{e}$ has a total load of at most $1$: $\left\|P_{\{v\}^{\prime}}(x_{e})\right\|_{1}=\frac{1}{\deg(v)}\left\|P(v)\right\|_{1}\leq 1$. Each non-split vertex of $V$ has load of 0.
@@
We combine demand states as we go up the decomposition tree $T$. We consider each cluster $S^{\prime}$ in $T$, and replace the demand states $P_{S_{1}^{\prime}},\ldots,P_{S_{r}^{\prime}}$ that are supported on its sub-clusters $S_{1}^{\prime},\ldots,S_{r}^{\prime}$ with a new demand state that is supported on $S^{\prime}$. This is done in Theorem 5.5 below. Note that our demand states are not necessarily valid demand states, since $P_{S^{\prime}_{i}}$ only stores the portion of the global demand that currently resides in $S^{\prime}_{i}$. The matching positive and negative parts of a commodity may lie in different clusters, so complete cancellation cannot occur within a single cluster. However, when summing over all demand states, these contributions do cancel out, and the total demand state remains valid. The demands states will, however, satisfy the following invariant with respect to $P$, the initial valid demand state we have fixed:
@@
It can be seen that $P_{\{v\}^{\prime}}$ satisfies the invariant with $\alpha=1$ for the cluster $\{v\}$.
@@
Let $S$ be a cluster, and let $S_{1},\ldots,S_{t}$ be its sub-clusters (i.e., $\{S_{1},\ldots S_{t}\}=\{L_{1},\ldots,L_{\ell}\}\cup\{R_{1},\ldots,R_{r}\}$), computed by Theorems 4.1 and 4.9 with parameter $0<\uptau\leq 1$. For each $1\leq i\leq t$, assume that $P_{S^{\prime}_{i}}$ is a demand state satisfying Invariant 5.4 for $S_{i}$ with load $\alpha\geq 1$. Let $P^{\mathrm{before}}_{S^{\prime}}=\sum_{i=1}^{t}P_{S^{\prime}_{i}}$ be their sum. Then, we can replace $P^{\mathrm{before}}_{S^{\prime}}$ with a demand state $P^{\mathrm{after}}_{S^{\prime}}$ which satisfies the following:
@@
    $G^{\prime}[S^{\prime}]$ $\Omega\!\left(\frac{\uptau}{\alpha\log n}\right)$-respects $P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}$. Moreover, $P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}$ is a valid demand state.
@@
Result (2) of Theorem 5.5 means that if we charge the edges $e\in E(U^{\prime},W^{\prime})\cap G^{\prime}[S^{\prime}]$ to cover the demand of $P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}$ across the cut $(U^{\prime},W^{\prime})$ by
$\operatorname{charge}_{S}(e)\mathrel{\mathop{:}}=\frac{\operatorname{dem}_{\left(P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}\right)}(U^{\prime},W^{\prime})}{\operatorname{cap}\!\left(E(U^{\prime},W^{\prime})\cap G^{\prime}[S^{\prime}]\right)},$
then $\operatorname{charge}_{S}(e)=O\!\left(\frac{\alpha\log n}{\uptau}\right)$. Note that the charged edges are in the subdivision graph.
In other words, Theorem 5.5 shows that the demand states supported on the sub-clusters $S_{1},\ldots,S_{t}$ can be merged into a single demand state on $S$, while only slightly increasing the load (by a factor of $1+O(\uptau)$) on the boundary edges of $S$. The cost of this operation is that we add $O\left(\frac{\alpha\log n}{\uptau}\right)$ to the charge of each edge $e\in E(U^{\prime},W^{\prime})\cap G^{\prime}[S^{\prime}]$. The proof of this theorem is given in Section 5.1.
@@
We prove that the conditions of Corollary 5.3 are satisfied. Let $P$, $(U,W)$, and $(U^{\prime},W^{\prime})$ be as in the discussion following Corollary 5.3. Set $\uptau=\frac{1}{\log n}$. We process $P$ in a bottom-up manner: First we convert it to a set of demand states $P_{\{v\}^{\prime}}$ as described above. Then, for each cluster in $T$, we apply Theorem 5.5 and Remark 5.6, accumulating the charges incurred at each application of the theorem. Let $\operatorname{charge}_{S}(e)$ denote the charge assigned to an edge $e\in E(U^{\prime},W^{\prime})\cap G^{\prime}[S^{\prime}]$ when Theorem 5.5 is applied to cluster $S$ as in Remark 5.6.
After applying the theorem for the final cluster $V$, the resulting demand state $P_{V^{\prime}}$ is actually the zero demand state on $V^{\prime}$. This follows from Invariant 5.4(1), since the cluster $V$ has no boundary edges. We now explain why the accumulated charges are enough to cover $\operatorname{dem}_{P}(U,W)$.
Denote by $\mathcal{Q}$ the set of current “active” clusters, such that initially $\mathcal{Q}=\mathcal{Q}_{0}\mathrel{\mathop{:}}=\{\{v\}\mid v\in V\}$. Then in each iteration we remove $S_{1},\ldots,S_{t}$ from $\mathcal{Q}$ and add $S$. Let $P_{\mathcal{Q}}$ be the sum of all demand states $P_{S^{\prime}}$, for $S\in\mathcal{Q}$.
@@
By induction on the applications of Theorem 5.5, for each $S\in\mathcal{Q}$, $P_{S^{\prime}}$ satisfies Invariant 5.4. Therefore, for each $k\in\mathcal{K}$,
@@
The following claim, whose proof is deferred to Appendix A, shows that $\operatorname{dem}_{P}(U,W)$ and $\operatorname{dem}_{P_{\mathcal{Q}_{0}}}(U^{\prime},W^{\prime})$ are the same up to some minor error that arises due to split nodes on the cut $(U^{\prime},W^{\prime})$ of $G^{\prime}$.
@@
$\operatorname{dem}_{P}(U,W)\leq\operatorname{dem}_{P_{\mathcal{Q}_{0}}}(U^{\prime},W^{\prime})+\operatorname{cap}_{G}(U,W)$.
Consider a replacement of $\mathcal{Q}$ by $\tilde{\mathcal{Q}}=\mathcal{Q}\cup\{S\}\setminus\{S_{1},S_{2},\ldots,S_{t}\}$ when we apply Theorem 5.5. Recall that $P^{\mathrm{before}}_{S^{\prime}}=\sum_{i=1}^{t}P_{S^{\prime}_{i}}$ is replaced by $P^{\mathrm{after}}_{S^{\prime}}$. Let $P^{\mathrm{before}}_{\tilde{\mathcal{Q}}}\mathrel{\mathop{:}}=P_{\mathcal{Q}}$ be the sum of active demand states before the change and let $P^{\mathrm{after}}_{\tilde{\mathcal{Q}}}\mathrel{\mathop{:}}=P_{\mathcal{Q}}-P^{\mathrm{before}}_{S^{\prime}}+P^{\mathrm{after}}_{S^{\prime}}$ be the sum after the change. We have $\operatorname{dem}_{\left(P^{\mathrm{before}}_{\tilde{\mathcal{Q}}}-P^{\mathrm{after}}_{\tilde{\mathcal{Q}}}\right)}(U^{\prime},W^{\prime})=\operatorname{dem}_{\left(P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}\right)}(U^{\prime},W^{\prime})$. This means that after accumulating all the charges in the applications of Theorem 5.5 on all clusters $S\in T$, the total charge is
$\sum_{e\in E(U^{\prime},W^{\prime})}{\operatorname{charge}(e)}$ $=\sum_{e\in E(U^{\prime},W^{\prime})}\sum_{S\in T}{\operatorname{charge}_{S}(e)}=\sum_{S\in T}\sum_{e\in E(U^{\prime},W^{\prime})\cap G^{\prime}[S^{\prime}]}{\operatorname{charge}_{S}(e)}$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu1\mkern-1.5mu)}}{{\geq}}\sum_{S\in T}{\operatorname{dem}_{\left(P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}\right)}(U^{\prime},W^{\prime})}=\sum_{S\in T}{\operatorname{dem}_{\left(P^{\mathrm{before}}_{\tilde{\mathcal{Q}}}-P^{\mathrm{after}}_{\tilde{\mathcal{Q}}}\right)}(U^{\prime},W^{\prime})}$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu2\mkern-1.5mu)}}{{\geq}}\operatorname{dem}_{\left(P_{\mathcal{Q}_{0}}-P_{V^{\prime}}\right)}(U^{\prime},W^{\prime})=\operatorname{dem}_{P_{\mathcal{Q}_{0}}}(U^{\prime},W^{\prime})\geq\operatorname{dem}_{P}(U,W)-\operatorname{cap}_{G}(U,W)~,$
where $\operatorname{charge}(e)$ is the total charge incurred on $e$, Inequality $(1)$ follows from Remark 5.6 and Inequality (2) uses Fact 3.16 on a telescoping sum. Recall that $P_{V^{\prime}}$ is the final demand state which is zero. By charging an additional unit to each cut edge, we get $\operatorname{dem}_{P}(U,W)\leq\sum_{e\in E(U^{\prime},W^{\prime})}{\operatorname{charge}(e)}$.
To conclude, we bound the total charge on each edge as follows. The height of the decomposition tree is $O(\log n)$ and $\uptau=\frac{1}{\log n}$, which means that $\alpha$ is a constant for each cluster. Therefore, by Theorem 5.5 and Remark 5.6, each edge is charged $O(\log^{2}n)$ for each cluster containing it, for a total charge of $O(\log^{3}n)$ per edge, which gives $\operatorname{dem}_{P}(U,W)=O(\log^{3}n)\cdot\operatorname{cap}_{G^{\prime}}(U^{\prime},W^{\prime})=O(\log^{3}n)\cdot\operatorname{cap}_{G}(U,W)$, as we wanted. Thus, the conditions of Corollary 5.3 are satisfied, and Theorem 2.9 follows. ∎
@@
The goal of this section is to prove Theorem 5.5. Recall the notations of this theorem: $S\in T$ is a cluster, and let $S_{1},\ldots,S_{t}$ be its sub-clusters (i.e., $\{S_{1},\ldots S_{t}\}=\{L_{1},\ldots,L_{\ell}\}\cup\{R_{1},\ldots,R_{r}\}$), computed by Theorems 4.1 and 4.9 with parameter $0<\uptau\leq 1$. For $i=1\ldots t$, $P_{S^{\prime}_{i}}$ is a demand state satisfying Invariant 5.4 for $S_{i}$ with a given load $\alpha\geq 1$.
Let $P^{\mathrm{before}}_{S^{\prime}}\mathrel{\mathop{:}}=\sum_{i=1}^{t}P_{S^{\prime}_{i}}$ be the sum of the demand states on the sub-clusters. By Invariant 5.4(1), all the non-zero demands in $P^{\mathrm{before}}_{S^{\prime}}$ are residing at the nodes in $X_{F}\cup X_{Y}\cup X_{B}$ of $S^{\prime}$ (these are the boundary edges of $\{S_{i}\}_{i=1}^{t}$). Given these demand vectors, we wish to “solve” some of these demands by updating the demand states such that a positive demand and a negative demand of the same commodity are moved to the same vertex, and cancel out. Then we want to “send” the remaining demands to $X_{B}$ (the boundary of $S$). In other words, we replace $P^{\mathrm{before}}_{S^{\prime}}$ by a demand state $P^{\mathrm{after}}_{S^{\prime}}$, where the demands of $P^{\mathrm{after}}_{S^{\prime}}$ lie on $X_{B}$ (i.e., satisfying Invariant 5.4 for $S$), and we show that $G^{\prime}[S^{\prime}]$ $\Omega(\frac{\uptau}{\alpha\log n})$-respects $P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}$.
We give a sketch of the argument via the following sequence of demand state updates:
@@
    Denote $P^{(1)}_{S^{\prime}}\mathrel{\mathop{:}}=P^{\mathrm{before}}_{S^{\prime}}$. First, we deal with demands that originated inside $L$, i.e., with the demand states $P^{(1)}_{L^{\prime}}=\sum_{i}{P_{L^{\prime}_{i}}}$, where $L=\bigcup_{i}{L_{i}}$. We move the demands of $P^{(1)}_{L^{\prime}}$ that reside on $X_{Y}\cap L^{\prime}$ to $X_{F}$, using the flow guaranteed by Property 2 of Theorem 4.9. This flow is routable in the graph with congestion $O(\frac{\alpha}{\uptau})$ and in particular, by Fact 2.7, is $\Omega(\frac{\uptau}{\alpha})$-respected by the graph. At this point, nodes in $X_{F}$ have $O(\frac{\alpha}{\uptau})$ load.
@@
    Then, we use a scaled all-to-all demand matrix on $X_{F}$ to update the demand state and spread the demands uniformly on $X_{F}$. By Theorem 4.1 we know that this problem is $\Omega(\frac{\uptau}{\alpha\log n})$-respected by the graph. This means that positive and negative demands of the same commodity cancel out, and we are only left with demand that has to enter or leave the cluster $L$. Because $L$ is a node in the hierarchical decomposition tree $T$ and $P$ is respected by $T$, the total amount of leftover demand is at most $\operatorname{cap}_{G}(L,V\setminus L)$.
@@
    Then, we apply the reverse demand transformations from step (1) to move the demands back to $X_{Y}$. [^14] In fact, the demands are moved to only some edges of $Y$ (see the proof for details).
@@
    Finally, we take all the demands that are currently on $X_{Y}$ (either from $P_{R^{\prime}}\mathrel{\mathop{:}}=P^{(1)}_{S^{\prime}}-P^{(1)}_{L^{\prime}}$ or the resulting updated demands), and use the flow ensured by Property 1 of Theorem 4.9 to route them to $X_{B}$. This flow is routable in the graph with congestion $O(\alpha)$ and, by Fact 2.7, is $\Omega(\frac{1}{\alpha})$-respected by the graph.
In all of these steps, we show that $G^{\prime}[S^{\prime}]$ respects the updates. We do this in two different ways: whenever the update is implemented by an explicit flow (as in steps 11,33 and $4$), we use the fact that the flow is routable with low congestion, and therefore, by Fact 2.7, the update is respected by $G^{\prime}[S^{\prime}]$. In the step that relies on uniform spreading on $X_{F}$ (step $2$), we instead use the expansion of $X_{F}$ to claim that the update is respected by $G^{\prime}[S^{\prime}]$. We formalize these arguments in the lemmas below.
Let $P^{(1)}_{L^{\prime}}\mathrel{\mathop{:}}=\sum_{i=1}^{\ell}P_{L^{\prime}_{i}}$ be the sum of the demand states on the sub-clusters within $L$. The following lemma captures steps $1$ and $2$ in the above process.
@@
    $G^{\prime}[S^{\prime}]$ $\Omega\!\left(\frac{\uptau}{\alpha\log n}\right)$-respects $P^{(1)}_{L^{\prime}}-P^{(3)}_{L^{\prime}}$. Moreover, $P^{(1)}_{L^{\prime}}-P^{(3)}_{L^{\prime}}$ is a valid demand state.
@@
Initially, Invariant 5.4(3) gives that every boundary split-node $x_{b}\in X_{B}$ in the graph $G^{\prime}[S^{\prime}]$ has load at most $\alpha$ in $P^{(1)}_{S^{\prime}}$, while vertices of $(X_{F}\cup X_{Y})\setminus X_{B}$ may have load at most $2\alpha$ (each of these vertices is a boundary node of two of the sub-clusters $\{S_{i}\}_{i=1}^{t}$, so two sub-clusters may have placed a load of $\alpha$ on it). That is, $\sum_{k\in\mathcal{K}}{|P^{(1)}_{S^{\prime}}(x_{b},k)|}\leq\alpha$ for boundary nodes $x_{b}\in X_{B}$, and $2\alpha$ for nodes in $(X_{F}\cup X_{Y})\setminus X_{B}$. Due to Invariant 5.4(1), $P^{(1)}_{L^{\prime}}$ is supported only on $L^{\prime}$.